# Labels Matter More Than Models: Quantifying the Benefit of Supervised Time Series Anomaly Detection

### v1

@@ Abstract
Time series anomaly detection (TSAD) is a critical data mining task often constrained by label scarcity. Consequently, current research predominantly focuses on Unsupervised Time-series Anomaly Detection (UTAD), relying on complex architectures to model normal data distributions. However, this approach often overlooks the significant performance gains available from limited anomaly labels achievable in practical scenarios. This paper challenges the premise that architectural complexity is the optimal path for TSAD. We conduct the first methodical comparison between supervised and unsupervised paradigms and introduce STAND, a streamlined supervised baseline. Extensive experiments on five public datasets demonstrate that: (1) Labels matter more than models: under a limited labeling budget, simple supervised models significantly outperform complex state-of-the-art unsupervised methods; (2) Supervision yields higher returns: the performance gain from minimal supervision far exceeds that from architectural innovations; and (3) Practicality: STAND exhibits superior prediction consistency and anomaly localization compared to unsupervised counterparts. These findings advocate for a data-centric shift in TSAD research, emphasizing label utilization over purely algorithmic complexity. The code is publicly available at https://github.com/EmorZz1G/STAND.
@@ I Introduction
Time series anomaly detection (TSAD) is a crucial and challenging task in time series data mining, with broad applications in fields such as industrial system monitoring, cybersecurity, and health surveillance [adamembls, cce, simad, TSB-UAD].
Due to the scarcity of anomaly samples and the high cost of labeling in TSAD, unsupervised time series anomaly detection (UTAD) methods have garnered significant attention in recent years [bls, simad, catch, M2N2]. Typically, unsupervised methods assume that the training time series data primarily consists of normal samples. They detect anomalies by learning the patterns of these normal samples, and this approach currently represents the mainstream of research. Because these methods introduce the prior assumption that the training data is composed entirely of normal samples, they are also referred to as semi-supervised techniques in some literature [TSB-UAD, catch, ppt]. Moving forward, we will denote this specific category as Unsupervised Time series Anomaly Detection Type II (UTAD-II), as illustrated in Figure 1(b).
Conversely, earlier studies trained and detected anomalies using unlabeled datasets that might contain anomaly samples, leading to their original designation as unsupervised techniques [TSB-UAD, pca]. We will refer to this category as Unsupervised Time series Anomaly Detection Type I (UTAD-I), as shown in Figure 1(a).
@@
In contrast, Supervised Time series Anomaly Detection (STAD) methods utilize anomaly sample labels during the training phase to optimize performance (as shown in Figure 1(c)), but such methods have received limited research attention [STSAD, ppt, label2].
@@
However, in practical applications, a small amount of anomaly sample labels can be acquired, and the purely unsupervised assumption often struggles to hold true [label1]. Existing studies largely overlook the significance of utilizing these few labels in time series anomaly detection, focusing instead primarily on improvements to model architecture.
@@
To address these critical questions, we propose a simple Supervised Time-series ANomaly Detection (STAND) baseline method. Furthermore, we present the first dedicated comparison between supervised and unsupervised methods to date. Through extensive experimentation, we derive the following key findings and contributions:
@@
    Labels Matter More Than Models: Under a limited labeling budget, the simple STAD method, STAND, significantly outperforms existing complex unsupervised time series anomaly detection methods. As illustrated in Figure 2, the STAD category consistently yields superior performance compared to both UTAD-I and UTAD-II.
@@
    The Benefit of Supervision: Given a limited labeling budget, the performance improvement gained from supervision far outweighs the incremental gains from model architecture improvements.
@@
    Readability & Practicality: Existing Unsupervised Time series Anomaly Detection (UTAD) methods often fail to clearly delineate or anchor anomaly segments; in contrast, STAD methods successfully detect the majority of anomalies while maintaining consistent predictions, demonstrating stronger practical applicability.
@@
    The First Supervised/Unsupervised TSAD Benchmark and Library: We present the first work to incorporate supervised methods into TSAD evaluation landscape. We established a complete evaluation pipeline and open-source our code [^1] STAND: https://github.com/EmorZz1G/STAND.
@@ II-A Algorithm-Centric Time Series Anomaly Detection
Early unsupervised Time Series Anomaly Detection (UTAD) methods were primarily based on classical machine learning techniques, such as Principal Component Analysis (PCA) [pca], Local Outlier Factor (LOF) [ml_lof2000], and Isolation Forest (IForest) [ml_IsolationForest2008]. These methods transform time series data into static feature vectors using techniques like sliding window segmentation and statistical feature extraction, which are then fed into traditional anomaly detection algorithms. However, these approaches heavily depend on the quality of hand-crafted features, and their model structures are often insufficiently adapted to the dynamic nature of time series data.
@@
Path I: Focusing on Model Architecture Improvement to Enhance Time Series Feature Extraction Capabilities. This includes methods based on improvements to classical neural networks [tcn_ed, usad, M2N2]; Transformer-based approaches [anom_trans], which leverage attention mechanisms to improve long-range dependency modeling; Mamba-based methods [s6_MemMambaAD, s6_wangMambaEffectiveTime2024, s6_behrouzChimeraEffectivelyModeling2024], which enhance the efficiency of processing long time sequences through state space models; and Large Language Model (LLM)-based methods [gpt2adpt, llm_AnomalyLLM2024], which utilize the general knowledge learned by LLMs for anomaly recognition.
Path II: Focusing on Establishing Anomaly Priors to Optimize Anomaly Determination. Examples include networks based on one-class classification and Random Forest principles [deep_svdd, deep_if], which introduce the concept of data centricity to identify anomalies; methods based on dissimilarity learning [patchad, simads, simad, dcdetector], which employ proxy strategies to learn the similarity distribution between normal and anomalous instances; and methods based on time-frequency information learning [TFAD, dual-tf, NPSR], which transform time series data into the frequency domain to capture multi-granularity anomalies.
@@ II-B Data-Centric Time Series Anomaly Detection
Unlike the goal of algorithm-centric research, data-centric time series anomaly detection research revolves around optimizing the data layer to compensate for the deficiency of labels in unsupervised settings. However, due to the low prevalence of anomaly events and the high cost of labeling in most practical scenarios, studies that directly leverage high-quality anomaly labels to optimize models remain scarce [STSAD, sv_landauerTimeSeriesAnalysis2018, sv_belayUnsupervisedAnomalyDetection2023]. Based on this constraint, past data-centric research on time series anomaly detection has mainly explored two optimization paths:
Path I: Pseudo-Sample-Based Data Augmentation Strategies to Expand Data Value with Limited Information. The core of this path is to substitute real labels with data transformations, creating pseudo-anomaly samples or task-oriented samples to provide learning signals for the model. This is implemented in two ways: First, unsupervised pseudo-sample generation, where normal samples are modified into anomaly samples in a label-free setting by simulating anomaly mechanisms (e.g., anomaly injection, segment shuffling, etc.). This guides the model to explore anomaly patterns and enhance detection capability [couta, CutAddPaste, trs_jeongAnomalyBERTSelfSupervisedTransformer2023]. Second, construction of task-guided samples, where, for instance, noisy time series data are treated as boundary samples between normal and anomalous states, or similar samples are generated via operations like Mixup, interpolation, or inversion to assist with model initialization [adamembls, amsl, simad, s6_zengCMambaChannelCorrelation2024].
Path II: New Sample Generation Strategies Based on Distribution Modeling, Guiding Pattern Learning through Data Generation. This path utilizes generative models (e.g., GANs, VAEs, Diffusion Models) to learn the distribution characteristics of normal time series data, thereby generating new samples that conform to realistic patterns and reinforcing the model’s perception of normal modes [GenIAS, app_liDiffTADDenoisingDiffusion2024a, gan_duGANBasedAnomalyDetection2021]. The generated samples can directly augment the training set, mitigating data scarcity. Simultaneously, an anomaly determination baseline is established through the detection of deviations from the learned normal distribution, forming a closed loop between data generation and anomaly detection.
@@ III-A Definitions
The multivariate time series sample is represented as $\mathbf{X}=[x_{1},x_{2},\dots,x_{T}]^{\top}\in\mathbb{R}^{T\times C}$, where $T$ is the time series length, $C$ denotes the number of variables, and $x_{t}\in\mathbb{R}^{C}$ is the multivariate observation at time $t$. The corresponding time-step labels are $\mathbf{y}=[y_{1},y_{2},\dots,y_{T}]^{\top}$, where $y_{t}=0$ indicates a normal time step and $y_{t}=1$ indicates an anomalous time step. The objective of anomaly detection is to learn a mapping function $f:\mathbb{R}^{T}\rightarrow\mathbb{R}$, which outputs the anomaly score for each time step, $\mathbf{S}=f(\mathbf{X})$. A preset threshold $\tau$ is then used for class determination: if $s_{t}>\tau$, time step $t$ is classified as anomalous; otherwise, it is normal.
@@ III-B Proposed Method
In this section, we introduce our proposed Supervised Time-series ANomaly Detection (STAND) framework. Our core hypothesis is that even a simple model, when guided by a small amount of supervision, can outperform complex unsupervised architectures. Consequently, STAND is intentionally designed with a simple yet effective architecture to serve as a strong baseline and empirically validate the significant impact of labels. The framework treats time-series anomaly detection as a sequence-to-sequence binary classification task at the timestamp level. As illustrated in Figure 3, STAND is comprised of three main components: a Feature Embedding, a Temporal Encoding Module, and an Anomaly Scoring Module.
@@ III-B1 Model Architecture
As defined in our implementation, this module is a simple Multi-Layer Perceptron (MLP) with two linear layers, each followed by a Gaussian Error Linear Unit (GELU) activation function [gelu] and Layer Normalization [layernorm]. The process for a single timestamp $x_{t}$ can be formulated as:
@@
Temporal Encoding Module. To capture the temporal dependencies inherent in time-series data, the sequence of embeddings $\mathbf{H}^{e}$ is passed to a Temporal Encoding Module. The choice of encoder is critical for modeling the contextual information that distinguishes an anomaly from normal behavior. We utilize a Long Short-Term Memory (LSTM) network [lstm], a proven and effective recurrent neural network (RNN) architecture for sequential data. LSTMs use a gating mechanism (input, forget, and output gates) to regulate information flow, enabling them to capture long-range dependencies while mitigating the vanishing gradient problem.
@@ III-B2 Training Objective
where $s_{t}$ is the output logit from the model for timestamp $t$, $y_{t}\in\{0,1\}$ is the corresponding ground-truth label, and $\sigma(\cdot)$ is the sigmoid function. The total loss is averaged over all samples in a mini-batch. The model parameters are updated via backpropagation using a standard gradient-based optimizer such as Adam [adam]. The complete training and inference procedures are summarized in Algorithm 1.
@@ IV Theoretical Analysis
To theoretically substantiate the efficacy of the proposed STAND framework, we provide a rigorous examination of its computational efficiency and optimization stability. This analysis focuses on the asymptotic complexity relative to input dimensions and establishes the convergence properties of the optimization objective.
IV-A Computational Complexity Analysis
The computational efficiency of STAND is a direct consequence of its streamlined architecture, which eschews the quadratic complexity often associated with canonical attention mechanisms.
@@ IV-A Computational Complexity Analysis
This formulation highlights a critical theoretical advantage: the complexity scales linearly with the sequence length $T$. In stark contrast, state-of-the-art unsupervised methods relying on self-attention mechanisms, such as Anomaly Transformer, exhibit a quadratic complexity $\mathcal{T}_{\text{Trans}}$:
@@ IV-B Convergence Analysis
$J(\theta)=-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left[y_{i,t}\log(\sigma(s_{i,t}))+(1-y_{i,t})\log(1-\sigma(s_{i,t}))\right]$ (14)
@@
While global optimality is not theoretically guaranteed due to the non-convex nature of the recurrent architecture, Eq. (14) rigorously confirms that STAND converges to a critical point (a local minimum or saddle point). The supervised nature of the loss landscape provides steeper gradients around anomalies compared to reconstruction-based unsupervised losses, thereby empirically accelerating this convergence rate.

### v2

# Labels Matter More Than Models: Rethinking the Unsupervised Paradigm in Time Series Anomaly Detection

@@ Abstract
Time series anomaly detection (TSAD) is a critical data mining task often constrained by label scarcity. Consequently, current research predominantly focuses on Unsupervised Time-series Anomaly Detection (UTAD), relying on increasingly complex architectures to model normal data distributions. However, this algorithm-centric trend often overlooks the significant performance gains achievable from limited anomaly labels available in practical scenarios. This paper challenges the premise that algorithmic complexity is the optimal path for TSAD. Instead of proposing another intricate unsupervised model, we present a comprehensive benchmark and empirical study to rigorously compare supervised and unsupervised paradigms. To isolate the value of labels, we introduce STAND, a deliberately minimalist supervised baseline. Extensive experiments on five public datasets demonstrate that: (1) Labels matter more than models: under a limited labeling budget, simple supervised models significantly outperform complex state-of-the-art unsupervised methods; (2) Supervision yields higher returns: the performance gain from minimal supervision far exceeds the incremental gains from architectural innovations; and (3) Practicality: STAND exhibits superior prediction consistency and anomaly localization compared to unsupervised counterparts. These findings advocate for a paradigm shift in TSAD research, urging the community to prioritize data-centric label utilization over purely algorithmic complexity. The code and benchmark are publicly available at https://github.com/EmorZz1G/STAND.
@@ I Introduction
In the era of big data, the proliferation of Internet of Things (IoT) devices and industrial sensors has generated massive volumes of streaming data [8, 22, 13, 10]. Consequently, Time Series Anomaly Detection (TSAD) has emerged as a crucial and challenging task in big data analytics, with broad applications in industrial system monitoring, cybersecurity, and health surveillance [50, 49, 51, 28, 46].
As the volume and velocity of time series data explode, acquiring high-quality anomaly labels becomes prohibitively expensive. Due to this severe label scarcity within massive datasets, unsupervised time series anomaly detection (UTAD) methods have garnered significant attention in recent years [48, 51, 41, 16]. Typically, unsupervised methods assume that the training time series data primarily consists of normal samples. They detect anomalies by learning the patterns of these normal samples, and this approach currently represents the mainstream of research. Because these methods introduce the prior assumption that the training data is composed entirely of normal samples, they are also referred to as semi-supervised techniques in some literature [28, 41, 26]. Moving forward, we will denote this specific category as Unsupervised Time series Anomaly Detection Type II (UTAD-II), as illustrated in Figure 1(b).
Conversely, earlier studies trained and detected anomalies using unlabeled datasets that might contain anomaly samples, leading to their original designation as unsupervised techniques [28, 31]. We will refer to this category as Unsupervised Time series Anomaly Detection Type I (UTAD-I), as shown in Figure 1(a).
@@
In contrast, Supervised Time series Anomaly Detection (STAD) methods utilize anomaly sample labels during the training phase to optimize performance (as shown in Figure 1(c)), but such methods have received limited research attention [15, 26, 56].
@@
However, in practical applications, a small amount of anomaly sample labels can be acquired, and the purely unsupervised assumption often struggles to hold true [54]. Existing studies largely overlook the significance of utilizing these few labels in time series anomaly detection, focusing instead primarily on improvements to model architecture.
@@
To address these critical questions, rather than pursuing marginal gains through architectural complexity, we propose a deliberately minimalist Supervised Time-series ANomaly Detection (STAND) baseline. STAND is intentionally stripped of overly complex mechanisms to serve as a pure conceptual probe, explicitly isolating and quantifying the true impact of supervisory signals. Furthermore, we establish a comprehensive unified benchmark bridging the gap between supervised and unsupervised TSAD evaluation. Through extensive empirical and theoretical analysis, we derive the following key findings and contributions:
@@
    Paradigm Shift via Empirical Evidence: We demonstrate that under a limited labeling budget, the simple STAD method (STAND) significantly outperforms existing highly complex unsupervised TSAD architectures. As illustrated in Figure 2, the STAD category consistently yields superior performance, proving our core thesis that ”Labels Matter More Than Models.”
@@
    Quantifying the Supervisory Gain: We theoretically and empirically validate that the performance improvement gained from minimal supervision (e.g., utilizing as little as 10% of the data) far outweighs the incremental gains typically achieved through unsupervised model architecture improvements.
@@
    Readability & Practicality in Real-world Scenarios: We reveal that existing UTAD methods often fail to clearly delineate anomaly segments and lack prediction consistency. In contrast, STAD methods successfully anchor anomalies while maintaining robust structural consistency, demonstrating significantly stronger practical applicability even under label noise.
@@
    A Unified Evaluation Benchmark and Library: Moving beyond mere model evaluation, we present a robust benchmark framework and an open-source library  [^1] https://github.com/EmorZz1G/STAND that successfully incorporates supervised methods into the TSAD evaluation landscape alongside unsupervised baselines, fostering future data-centric TSAD research.
@@ II-A Algorithm-Centric Time Series Anomaly Detection
Early unsupervised Time Series Anomaly Detection (UTAD) methods were primarily based on classical machine learning techniques, such as Principal Component Analysis (PCA) [31], Local Outlier Factor (LOF) [5], and Isolation Forest (IForest) [24]. These methods transform time series data into static feature vectors using techniques like sliding window segmentation and statistical feature extraction, which are then fed into traditional anomaly detection algorithms. However, these approaches heavily depend on the quality of hand-crafted features, and their model structures are often insufficiently adapted to the dynamic nature of time series data.
@@
Path I: Focusing on Model Architecture Improvement to Enhance Time Series Feature Extraction Capabilities. This includes methods based on improvements to classical neural networks [9, 1, 16]; Transformer-based approaches [44], which leverage attention mechanisms to improve long-range dependency modeling; Mamba-based methods [21, 39, 3], which enhance the efficiency of processing long time sequences through state space models; and Large Language Model (LLM)-based methods [55, 23], which utilize the general knowledge learned by LLMs for anomaly recognition.
Path II: Focusing on Establishing Anomaly Priors to Optimize Anomaly Determination. Examples include networks based on one-class classification and Random Forest principles [29, 42], which introduce the concept of data centricity to identify anomalies; methods based on dissimilarity learning [52, 53, 51, 45], which employ proxy strategies to learn the similarity distribution between normal and anomalous instances; and methods based on time-frequency information learning [47, 25, 18], which transform time series data into the frequency domain to capture multi-granularity anomalies.
@@ II-B Data-Centric Time Series Anomaly Detection
Unlike the goal of algorithm-centric research, data-centric time series anomaly detection research revolves around optimizing the data layer to compensate for the deficiency of labels in unsupervised settings. However, due to the low prevalence of anomaly events and the high cost of labeling in most practical scenarios, studies that directly leverage high-quality anomaly labels to optimize models remain scarce [15, 19, 4]. Based on this constraint, past data-centric research on time series anomaly detection has mainly explored two optimization paths:
Path I: Pseudo-Sample-Based Data Augmentation Strategies to Expand Data Value with Limited Information. The core of this path is to substitute real labels with data transformations, creating pseudo-anomaly samples or task-oriented samples to provide learning signals for the model. This is implemented in two ways: First, unsupervised pseudo-sample generation, where normal samples are modified into anomaly samples in a label-free setting by simulating anomaly mechanisms (e.g., anomaly injection, segment shuffling, etc.). This guides the model to explore anomaly patterns and enhance detection capability [43, 38, 14]. Second, construction of task-guided samples, where, for instance, noisy time series data are treated as boundary samples between normal and anomalous states, or similar samples are generated via operations like Mixup, interpolation, or inversion to assist with model initialization [50, 51].
Path II: New Sample Generation Strategies Based on Distribution Modeling, Guiding Pattern Learning through Data Generation. This path utilizes generative models (e.g., GANs, VAEs, Diffusion Models) to learn the distribution characteristics of normal time series data, thereby generating new samples that conform to realistic patterns and reinforcing the model’s perception of normal modes [7, 20]. The generated samples can directly augment the training set, mitigating data scarcity. Simultaneously, an anomaly determination baseline is established through the detection of deviations from the learned normal distribution, forming a closed loop between data generation and anomaly detection.
@@ III-A Definitions
The multivariate time series sample is represented as $\mathbf{X}=[x_{1},x_{2},\dots,x_{T}]^{\top}\in\mathbb{R}^{T\times C}$, where $T$ is the time series length, $C$ denotes the number of variables, and $x_{t}\in\mathbb{R}^{C}$ is the multivariate observation at time $t$. The corresponding time-step labels are $\mathbf{y}=[y_{1},y_{2},\dots,y_{T}]^{\top}$, where $y_{t}=0$ indicates a normal time step and $y_{t}=1$ indicates an anomalous time step. The objective of anomaly detection is to learn a mapping function $f:\mathbb{R}^{T\times C}\rightarrow\mathbb{R}^{T}$, which outputs the anomaly score sequence for each time step, $\mathbf{S}=f(\mathbf{X})$. A preset threshold $\tau$ is then used for class determination: if $s_{t}>\tau$, time step $t$ is classified as anomalous; otherwise, it is normal.
@@ III-B Proposed Method
In this section, we introduce our proposed Supervised Time-series ANomaly Detection (STAND) framework. Our core hypothesis is that even a rudimentary model, when guided by a small amount of supervision, can decisively outperform complex unsupervised architectures.
Consequently, STAND is intentionally designed with a minimalist and structurally straightforward architecture. By deliberately eschewing highly parameterized or overly intricate components (such as deep self-attention mechanisms or complex generative bottlenecks), we ensure that STAND serves as a pure proof-of-concept baseline. This minimalist design guarantees that the observed performance superiority stems entirely from the utilization of supervisory signals rather than architectural over-engineering, thereby empirically validating the significant impact of labels. The framework treats time-series anomaly detection as a sequence-to-sequence binary classification task at the timestamp level. As illustrated in Figure 3, STAND is comprised of three basic components: a Feature Embedding, a Temporal Encoding Module, and an Anomaly Scoring Module.
@@ III-B1 Model Architecture
As defined in our implementation, this module is a simple Multi-Layer Perceptron (MLP) with two linear layers, each followed by a Gaussian Error Linear Unit (GELU) activation function [11] and Layer Normalization [2]. The process for a single timestamp $x_{t}$ can be formulated as:
@@
Temporal Encoding Module. To capture the temporal dependencies inherent in time-series data, the sequence of embeddings $\mathbf{H}^{e}$ is passed to a Temporal Encoding Module. The choice of encoder is critical for modeling the contextual information that distinguishes an anomaly from normal behavior. We utilize a Long Short-Term Memory (LSTM) network [12], a proven and effective recurrent neural network (RNN) architecture for sequential data. LSTMs use a gating mechanism (input, forget, and output gates) to regulate information flow, enabling them to capture long-range dependencies while mitigating the vanishing gradient problem.
@@ III-B2 Training Objective
where $s_{t}$ is the output logit from the model for timestamp $t$, $y_{t}\in\{0,1\}$ is the corresponding ground-truth label, and $\sigma(\cdot)$ is the sigmoid function. The total loss is averaged over all samples in a mini-batch. The model parameters are updated via backpropagation using a standard gradient-based optimizer such as Adam [17]. The complete training and inference procedures are summarized in Algorithm 1.
@@ IV Theoretical Analysis
To theoretically substantiate the proposed STAND framework and our core motivation, we rigorously examine its scalability, optimization stability, and sample complexity.
IV-A Scalability Analysis
The highly scalable nature of STAND is rooted in its computational efficiency, a direct consequence of its streamlined architecture, which eschews the quadratic complexity often associated with canonical attention mechanisms.
@@ IV-A Scalability Analysis
This formulation highlights a critical theoretical advantage regarding scalability for big data applications: the computational complexity scales strictly linearly with the sequence length $T$, making STAND highly efficient for processing massive, long-term time series streams. In stark contrast, state-of-the-art unsupervised methods relying on self-attention mechanisms, such as Anomaly Transformer, exhibit a quadratic complexity $\mathcal{T}_{\text{Trans}}$:
@@
Meanwhile, the runtime memory required for storing hidden states for backpropagation scales strictly linearly:
$\mathcal{M}_{\text{run}}=\mathcal{O}(T\cdot d\cdot L)$ (14)
Similar to the computational cost, this linear memory dependence offers a crucial scalability advantage over Transformer-based models, which typically require $\mathcal{O}(T^{2})$ memory for attention matrices. This ensures that STAND maintains a manageable memory footprint even when processing massive data streams, further cementing its suitability for large-scale industrial deployments.
@@ IV-B Convergence Analysis
$\small J(\theta)=-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left[y_{i,t}\log(\sigma(s_{i,t}))+(1-y_{i,t})\log(1-\sigma(s_{i,t}))\right]$ (15)
@@
While this represents a standard descent argument and global optimality is not theoretically guaranteed due to the non-convex nature of the recurrent architecture, Eq. (14) rigorously confirms that STAND converges stably to a critical point. We include this analysis to draw a deliberate contrast: unlike state-of-the-art UTAD methods relying on adversarial training (e.g., GANs) or complex reconstruction bottlenecks—which are notoriously prone to mode collapse and training instability—our supervised objective maintains standard, reliable convergence properties. This theoretical stability translates directly into the empirical robustness and prediction consistency observed in our evaluations.
IV-C Sample Complexity and the Benefit of Supervision
To rigorously substantiate our empirical finding that labels matter more than models, we analyze the intrinsic learning difficulty of Unsupervised Time-series Anomaly Detection (UTAD) versus Supervised Time-series Anomaly Detection (STAD) from the perspective of Statistical Learning Theory [37, 30]. Specifically, we formulate their respective risk functions and compare the required sample complexities to achieve a bounded generalization error $\epsilon$.
Unsupervised Complexity (UTAD): UTAD methods fundamentally rely on modeling the intrinsic distribution of normal data within a high-dimensional feature space $\mathbb{R}^{C}$. Let $p^{*}(x)$ be the true probability density function of the normal data, and $\hat{p}(x)$ be the estimated density function. In this paradigm, identifying anomalies equates to non-parametric density estimation, where the learning objective is to minimize the $L_{\infty}$ risk:
$\mathcal{R}_{\text{UTAD}}(\hat{p})=\sup_{x\in\mathbb{R}^{C}}|\hat{p}(x)-p^{*}(x)|$ (23)
According to minimax lower bounds in non-parametric statistics [34, 32], estimating a density function that is $s$-times differentiable in $\mathbb{R}^{C}$ with an expected error bound $\mathbb{E}[\mathcal{R}_{\text{UTAD}}(\hat{p})]\leq\epsilon$ requires a sample size $N_{\text{UTAD}}$ satisfying:
$N_{\text{UTAD}}=\Omega\left(\epsilon^{-\frac{C}{s}}\right)$ (24)
This exponential dependence on the variable dimension $C$ manifests the curse of dimensionality. Consequently, as the number of time-series channels increases (e.g., $C=122$ in the WADI dataset), purely unsupervised methods demand an exponentially larger normal dataset to accurately map the data manifold. This theoretical bottleneck explains why current algorithm-centric UTAD methods, despite employing complex architectures, are highly susceptible to overfitting background noise and producing false positives.
Supervised Complexity (STAD): Conversely, STAD reframes the objective from high-dimensional density estimation to a discriminative binary classification task. Let $\mathcal{H}$ be the hypothesis class of the supervised model (e.g., the neural architecture of STAND) with a Vapnik-Chervonenkis (VC) dimension of $d_{\text{VC}}(\mathcal{H})$. For a specific classifier $h\in\mathcal{H}$ and label sequence $\mathbf{y}$, the expected risk $R(h)$ and empirical risk $\hat{R}(h)$ are defined as:
$R(h)=\mathbb{E}_{(x,y)\sim\mathcal{P}_{X,Y}}[\ell(h(x),y)]$ (25)
$\hat{R}(h)=\frac{1}{N_{\text{STAD}}}\sum_{i=1}^{N_{\text{STAD}}}\ell(h(x_{i}),y_{i})$ (26)
By leveraging the explicit labels, the algorithm learns a decision boundary via Empirical Risk Minimization. Under the agnostic Probably Approximately Correct (PAC) learning framework [36], the generalization gap is bounded with a probability of at least $1-\delta$ by:
$\sup_{h\in\mathcal{H}}|R(h)-\hat{R}(h)|\leq\mathcal{O}\left(\sqrt{\frac{d_{\text{VC}}(\mathcal{H})+\ln(1/\delta)}{N_{\text{STAD}}}}\right)$ (27)
To ensure the generalization error is strictly bounded by $\epsilon$, the required labeled sample size $N_{\text{STAD}}$ converges to:
$N_{\text{STAD}}=\mathcal{O}\left(\frac{d_{\text{VC}}(\mathcal{H})+\ln(1/\delta)}{\epsilon^{2}}\right)$ (28)
This bound reveals that the sample complexity of STAD grows only polynomially with respect to $1/\epsilon$ and linearly with the model’s capacity $d_{\text{VC}}(\mathcal{H})$, entirely circumventing the exponential penalty $\epsilon^{-C/s}$.
Remark: The theoretical divergence between Eq. (24) and Eq. (28) mathematically formalizes our core motivation: introducing supervisory signals collapses an exponentially hard distribution-matching problem into a polynomially solvable discrimination problem. This rigorously explains the phenomena observed in our experiments (Sec. V-D), where a simple supervised baseline trained on a drastically reduced data budget (e.g., 20% of the labeled data) can decisively outperform state-of-the-art unsupervised architectures attempting to learn the full input distribution.